OpenStudy (anonymous):

Can someone help me to solve the following equation system. y² =x+1 x² + y² =13

OpenStudy (anonymous):

x=-4 and x=3?

OpenStudy (anonymous):

x^2 + (x+1) = 13 x^2 + x -12 = 0 (x+4)(x-3) = 0 x= -4 , x = 3

OpenStudy (anonymous):

substitute the first equation y^2 = x + 1 into the second equation x^2 + x + 1 = 13 now you have a polynomial equation, just re arrange to x^2 + x + 1-13 = 0 and use the quadratic formula to solve

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

then make sure you solve for y